FOR A SUM OF 8000 TO DOUBLE ITSELF IN 12 YRS. AND 9 MOS., WHAT MUST BE THE RATE OF INTEREST COMPOUNDED QUARTERLY

Answer by lwsshak3(11628)

FOR A SUM OF 8000 TO DOUBLE ITSELF IN 12 YRS. AND 9 MOS., WHAT MUST BE THE RATE OF INTEREST COMPOUNDED QUARTERLY
**
compound interest formula:
A=P(1+r)^n, r=interest rate per period, n=number of periods, P=initial investment, A=amt after n-periods.
For given problem:
A/P=(1+r)^n
A/P=2
r=to find
12 yrs=4*12=48 quarters
9 months=3 quarters
n=48+3=51 quarters
equation:
(1+r)^n=A/P
(1+r)^51=2
raise both sides to (1/51)
1+r=2^(1/51)
r=2^(1/51)-1
r≈.0137/quarter
or
r≈.0547/yr
Rate of interest compounded quarterly≈5.47%